OpenStudy (anonymous):

f(x)=1-3^x the graph of f is reflected over the line y=x. what is the name of function of the new graph?

OpenStudy (saifoo.khan):

i think that will be a straight line through which this will be reflected.

OpenStudy (anonymous):

the inverse 0.o

OpenStudy (anonymous):

f¯¹(x)

OpenStudy (saifoo.khan):

i have no idea.

OpenStudy (anonymous):

y = 1-3^x x = 1-3^y then solve for y.. I.. don't know how to do that for this problem...

OpenStudy (anonymous):

you are going to have to use logarithms

OpenStudy (anonymous):

\[ \log_{3}\left| -x+3 \right| =\log_{3}3^y\]\[y=\log_{3}\left| -x+3 \right|\] I think I could be very very wrong lol

OpenStudy (anonymous):

saifoo its add maths funtion

OpenStudy (dumbcow):

\[y = 1 - 3^{x}\] switch variables \[x = 1-3^{y}\] solve for y \[3^{y} +x = 1\] \[3^{y} = 1-x\] \[\log 3^{y} = \log (1-x)\] \[y*\log 3 = \log(1-x)\] \[y = \frac{\log (1-x)}{\log 3}\]

OpenStudy (anonymous):

what is the name of function of the new graph? no name

OpenStudy (dumbcow):

\[f^{-1}(x)\]

OpenStudy (anonymous):

f(x)\[=1-3^x\] f(x)=y \[f(y)^{-1}=x\] y=1-3^x 1-y=3^x \[\log_{3}(1-y)=x \] \[f(y)^{-1}=x\] \[f(y)^{-1}\]=\[\log_{3}(1-y)\] if u insert x instead of y \[f(x)^{-1}\]=\[\log_{3}(1-x)\]